OpenStudy (anonymous):

Which two points are on the graph of F(e)=10-(the fraction) 2 over 5e ? Choose exactly two answers that are correct. (Points : 1) (0, –10) (0, 10) (4, 0) (25, 0)

OpenStudy (anonymous):

@Karon86

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

x is a variable. I am going to use x instead of e. e confuses me too much. F(x) = 10 - (2/5x) you would have to exclude A and B, because then if you plug in zero for e, you get an undefined value. The function doesn't exist at x=0.

OpenStudy (solomonzelman):

if the y-coordinate is 0, then solve for the x. 0 = 10 - (2/5x) -10 = -(2/5x) 10 = (2/5x) multiply times 5x on both sides 50x = 2 x=1/25

OpenStudy (solomonzelman):

Are you sure you have all the options correctly, perhaps D is ( 1/25 , 0 ) ?

OpenStudy (anonymous):

im sure. the 2/5 is a fraction though

OpenStudy (solomonzelman):

wait x is not the denominator, right?

OpenStudy (solomonzelman):

F(x) = 10 - (2/5)x like this?

OpenStudy (solomonzelman):

then we would need to solve for without excluding A and B.

OpenStudy (solomonzelman):

F(x) = 10 - (2/5)x if x=0 F(x) = 10 - (2/5)(0) F(x)=10

OpenStudy (anonymous):

hold on

OpenStudy (solomonzelman):

see when x=0, y=10, so B is on it

OpenStudy (anonymous):

and there is 2 ansers from the choices i put

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

are you there???

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@iambatman